OpenStudy (anonymous):

Find the product of (4x + 3y)(4x − 3y). 16x2 − 24xy + 9y2 16x2 − 9y2 16x2 + 24xy + 9y2 16x2 + 9y2

OpenStudy (anonymous):

\[16x^2-9y^2\]

OpenStudy (anonymous):

please explain

OpenStudy (anonymous):

4x*4x=16x^2 4x*-3y=-12xy 3y*4x=12xy 3y*-3y=-9y^2 Just add them together